"a stone is throwen in a vertically up wards directionwith the velocity of 5 m/s.if the acceleration of the stone is 10m/s in the downwords direction .what will be the height attained by the stone and how much time will it take t

Grade:9

1 Answers

Vaibhav
9 years ago
u=5m/s a= -10m/s^2 h=Max. Height
:v^2=U^2-2gH
0=25-20H
20H=25
:H=5/4m
 
Now,
v=u-gt
0=5-10t
10t=5
t=1/2s